See Hint Kasey has income of $20.00. She always consumes 5 ounces of cream cheese with one bagel. If the price of cream cheesd
Answer #1

Answer : Spending amount for 1 time consumption = (price of per ounce cream cheese * quantity of cream cheese) + (price of a bagel * quantity of bagel) = (0.5 * 5) + (1.50 * 1) = $4

Kasey's income = $20

So, with the income of $20 Kasey can consume (20 / 4) = 5 times

For 5 times the number of cream cheese consumption is (5 * 5) = 25 ounce of cream cheese

For 5 times the number of bagel consumption is (5 * 1) = 5 bagels

Therefore, by purchasing 5 bagels and 25 ounce of cream cheese Kasey will maximize the utility.